OpenStudy (anonymous):

Name at least 2 different ways you can solve the equation 5^(2x+1)=25

OpenStudy (lynfran):

exponential...and i think logs

OpenStudy (anonymous):

i have one method so far, which is solving by having equal bases is that right?

OpenStudy (lynfran):

5^(2x+1)=25 5^(2x+1)=5^2 so 2x+1=2 2x=2-1 x=1/2

OpenStudy (lynfran):

yes equal base ... same as exponent

OpenStudy (lynfran):

im logs also

OpenStudy (anonymous):

would you solve the problem the same way with logs? or do you solve it with different steps?

OpenStudy (lynfran):

\[\log _{25}(5)=(2x+1)\]